A card is dealt from a complete deck of fifty-two playing cards (no jokers). Use probability rules (when appropriate) to find the probability that the card is as stated. (Enter your answers as fractions.)

(a) a king and a spade


(b) a king or a spade


(c) not a king of spades
Answer #1

  

Probability of event A(card being king) = 4/52
Probabily of event B(card being spade) =13/52

a) Probability of event (A\bigcap \bigca B) ie a king and a spade = 1/52

b) As events A and B are dependent events( there is possibility of king of spade)

P(A or B) = P(A) +P(B) -P(A and B)

Answer= 4/52 + 13/52 -1/52 = 16/52

c) probability of not a king of spade= 1- probability of a king of a spade

= 1 - 1/52

= 51/ 52